A woman is 15 years older than her younger brother. Six years ago, her brother is one - half as old as his elder sister. How old are they now?
Answer by ankor@dixie-net.com(22740)

let w = the woman's present age
let b = the brother's age
:
write an equation for each statement
:
A woman is 15 years older than her younger brother.
w = b + 15
Six years ago, her brother is one - half as old as his elder sister.
b - 6 = .5(w - 6)
b - 6 = .5w - 3
b = .5w - 3 + 6
b = .5w + 3
Replace w with (b+15)
b = .5(b+15) + 3
b = .5b + 7.5 + 3
b = .5b + 10.5
b - .5b = 10.5
.5b = 10.5
multiply both sides by 2
b = 21 yrs is the brothers age
then
w = 21 + 15
w = 36 yrs is the woman's age
:
:
Check this in the statement:
"Six years ago, her brother is one - half as old as his elder sister."
21 - 6 = .5(36 - 6)
15 = .5(30)